# FELLE: Autoregressive Speech Synthesis with Token-Wise Coarse-to-Fine Flow Matching

### v1

@@ 4.2 Model Detail
Our model comprises 12 Transformer blocks, designed in line with the architectures of VALL-E (Chen et al., 2025) and MELLE (Meng et al., 2024) to ensure a fair comparison. Each block features 16 attention heads and a feed-forward layer with a dimensionality of 4,096. The decoder incorporates an embedding dimension of 1,024 and ReLU activation function. The input mel-spectrograms are transformed into the model’s embedding space using a three-layer fully connected network. HiFi-GAN vocoder [^1] https://huggingface.co/mechanicalsea/speecht5-tts (Kong et al., 2020) is used for audio reconstruction.
The downsampling operation retains even-indexed mel-spectrogram frames as coarse components. During upsampling, these components are expanded through zero-insertion at odd-indexed positions. Both coarse and fine flow-matching stages share identical backbone architectures comprising three residual blocks, each containing layer normalization, dual fully connected layers, and SiLU activation. The timestep embedding module combines sinusoidal positional encoding with two fully connected layers and SiLU activation. Key architectural differences emerge in the conditioning modules: the coarse stage uses single linear projections for language model outputs, and the fine stage incorporates additional layers to integrate auxiliary features like coarse-mel information.
@@
The model is trained for 2 million iterations on 8 NVIDIA V100 GPUs using the AdamW optimizer. Loss coefficients are configured with $\beta=0.1$ for the condition loss $\mathcal{L}_{\text{cond}}$ and $\sigma=0.01$ for the stop prediction loss $\mathcal{L}_{\text{stop}}$. The noise variance during training is configured at 0.1. CFG is implemented with dropout probability $p_{\text{drop}}=0.1$. For unconditional setting, we apply a mask of random length between 3 and 10 seconds to audio and apply random audio masking (3-10 seconds duration) on speech prompt. During the inference process, we simultaneously process two types of input: one with complete speech prompts and another with masked speech prompts, used as classifier-free guidance. The results are then combined using a weighting factor of $w=1.6$ to produce the final output. The flow-matching framework performs 3 function evaluations using Euler’s method to iteratively generate mel-spectrograms.
4.3 Evaluation
Evaluation Setting
@@ 4.3 Evaluation
Continuation: Using the text transcription and the initial 3 seconds of the utterance as a prompt, the model is tasked with seamlessly synthesizing the continuation of the speech.
Cross-sentence: Given a reference utterance and its transcription from the same speaker as the prompt, along with the text of the target utterance, the model is expected to synthesize the corresponding speech while preserving the speaker’s distinctive characteristics.
Evaluation Metric
To assess the naturalness, robustness, speaker similarity, and effectiveness of the proposed method, we adopt multiple subjective and objective metrics:
Word Error Rate (WER): We use two automated speech recognition (ASR) models, the Conformer-Transducer model [^2] https://huggingface.co/nvidia/stt_en_conformer_transducer_xlarge (Gulati et al., 2020) and the HuBERT-Large ASR model [^3] https://huggingface.co/facebook/hubert-large-ls960-ft (Hsu et al., 2021), to evaluate the robustness and intelligibility of the synthesized speech. By contrasting the transcriptions produced by these models with the matching ground truth, the WER is determined. In particular, the WER scores derived from the Conformer-Transducer and HuBERT-Large systems are indicated by WER-C and WER-H, respectively.
Speaker Similarity (SIM): WavLM-TDNN [^4] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification (Chen et al., 2022) is employed to extract speaker embeddings from the reference speech and the synthesized speech to assess the in-context learning capability of zero-shot TTS models. The cosine distance, which goes from -1 to 1, is used to measure how similar these embeddings are to one another. There are two assessment indicators taken into account: SIM-r measuring the similarity between synthesized speech and the reconstructed speech prompt, and SIM-o comparing synthesized speech with the original prompt.
Mean Opinion Score (MOS): With the rapid development of automatic MOS prediction technology (Cooper et al., 2022; Wang et al., 2024, 2023b; Liu et al., 2025), it is now possible to evaluate quality accurately and effectively. We use the RAMP+ model [^5] https://github.com/NKU-HLT/RAMP_MOS (Wang et al., 2023a) for speech quality assessment in Appendix A.1.
5 Results
Table 1: The objective performance comparison of the continuation and cross-sentence zero-shot speech synthesis tasks, with WER-C (%) and WER-H (%) as WER metrics. *The reproduction results are quoted from Han et al. (2024). Results reported in Meng et al. (2024) are used. Meng et al. (2024) provides results of MELLE across various setups, and we show the performance on the same dataset.
@@ 5 Results
5.1 Comparative Study
We compare FELLE with several speech synthesis models, including ELLA-V, VALL-E R, RALL-E, CLAM-TTS, VALL-E, and MELLE in Table 1. These models are strong baselines in the field of speech synthesis. These models represent a diverse range of approaches, encompassing both discrete representation-based and continuous representation-based methods, providing a valuable benchmark for evaluating our approach. Additionally, the performance of ground truth (GT) is included as a reference point to highlight the upper bounds of these metrics. Another GT, labeled as GT-mel, refers to the reconstruction from the mel-spectrogram with the vocoder.
@@ 5.1 Comparative Study
5.2 Ablation Study
Table 2: The ablation study from two perspectives: the prior distribution and the generation mechanism. The ✓ denotes methods used in our paper, while others refer to ablation methods. ‘Vanilla Prior’ denotes a standard Gaussian distribution as the prior, ‘HFM’ denotes holistic flow matching that generates complete mel-spectrogram frames through a unified flow-matching process, and ‘DFM’ represents decoupled flow matching where separate flow-matching processes independently generate low- and high-frequency components without cross-band condition.
@@ 5.2 Ablation Study
5.3 Parameter Analysis
Figure 3: The TTS performance on continuation task using the Euler solver across different NFEs. The x-axis in both plots represents NFE. The left plot demonstrates the effect of NFE on WER metrics (WER-C and WER-H). The right plot illustrates the impact of NFE on similarity (SIM-o and SIM-r). The results are averaged over multiple runs conducted on different checkpoints to ensure robustness and reliability.
@@ 5.3 Parameter Analysis
The impact of the number of function evaluations (NFE) on TTS performance is shown in Figure 3. In our experiments, we employ the Euler method for numerical integration, where each step corresponds to one function evaluation. Initially, WER decreases as NFE increases, reflecting the improved mapping between prior distributions and acoustic outputs achieved by flow matching. This enhancement leads to better intelligibility and clarity in the generated speech. However, beyond a certain threshold, WER begins to rise, likely due to overfitting or the introduction of excessive distortion caused by additional iterations. In contrast, the similarity metric consistently declines as NFE increases. This suggests that excessive refinement may result in over-smoothing or the loss of fine-grained details, ultimately reducing perceptual similarity. These observations highlight a critical trade-off in the use of flow matching: while a moderate number of NFE improves clarity and reduces WER, excessive NFE can degrade both intelligibility and similarity. Therefore, careful tuning of NFE is essential to strike an optimal balance between clarity and similarity.
Figure 4: The continuation performance of different CFG scales ranges from 1 to 3 with a stride of 0.2. The x-axis in both plots represents the CFG scale. The left plot shows WER metrics (WER-C, WER-H), and the right plot shows similarity metrics (SIM-o, SIM-r). A CFG scale of 1 indicates no CFG is applied.
CFG Scale
We also explore the impact of the C2F-FM network scale and the prior distribution on the performance of flow matching, with the results presented in Appendix A.2 and Appendix A.3, respectively.
6 Conclusion
7 Limitation and Ethics Statement
While our proposed speech synthesis model demonstrates promising performance in preliminary experiments, several limitations remain that need to be addressed. Firstly, the evaluation has been primarily conducted on a limited language scope, focusing mainly on English. This restricts our ability to assess the model’s generalization ability across a broader range of languages, particularly low-resource languages or dialects with complex phonological features. Secondly, the dataset used for evaluation is of moderate size, and further exploration is needed on larger datasets to better assess the model’s performance and robustness. Additionally, experiments have only been conducted on mel-spectrograms, and further investigation is required to determine whether our method can also perform well with other continuous representations.
FELLE is a research project focused on speech synthesis that preserves speaker identity, with no current plans for commercialization or public release. It has potential applications in education, entertainment, accessibility, and more, but its output quality depends on factors like input quality and background noise. While FELLE can mimic specific voices, it also poses risks, such as voice spoofing or impersonation. To address these, a protocol ensuring speaker consent and a detection model for synthesized speech is essential. FELLE commits to ethical AI practices, and any misuse can be reported through the Report Abuse Portal.
Appendix A Appendix
A.1 MOS
Table 3: The predicted MOS on the LibriSpeech test-clean dataset for continuation and cross-sentence scenarios.
System Continuation Cross-Sentence
Ground Truth $4.043_{\pm 0.3165}$ $4.043_{\pm 0.3165}$
VALL-E [Chen et al., 2025] $1.828_{\pm 0.2373}$ $1.965_{\pm 0.2713}$
MELLE [Meng et al., 2024] $\textbf{3.843}_{\pm 0.3847}$ $4.036_{\pm 0.2538}$
FELLE $3.836_{\pm 0.3882}$ $\textbf{4.157}_{\pm 0.1891}$
A.2 Scaling of the C2F-FM Network
The experimental results in Table 4 highlight the trends of WER and similarity as the scale of the C2F-FM network increases. The smallest model (3×512) demonstrates limited learning capacity, resulting in high WER and poor similarity scores, suggesting the limitations of model size on its capabilities. Further scaling to 3×1024 achieves a better balance, with both WER and similarity metrics showing notable improvements. However, when the model size is increased further to 6×1024 and 12×1024, the gains in WER become marginal, and similarity scores begin to degrade, indicating potential overfitting. This suggests that while larger models can capture more complex patterns, excessive scaling may harm generalization, particularly in maintaining high similarity.
@@ A.2 Scaling of the C2F-FM Network
Table 4: Comparison of performance with the scale of the C2F-FM Network on continuation and cross-sentence zero-shot speech synthesis tasks. The first column indicates the number of residual blocks and the width of fully-connected layers in each block (e.g., 3×512 denotes 3 residual block with 512-dimensional layers). Total parameters (# Params) represent the sum of trainable parameters in both coarse and fine flow-matching networks.
@@
In conclusion, simply increasing the scale of the C2F-FM network is not sufficient to achieve optimal performance. Future work should focus on developing more efficient architectures and training strategies to improve both WER and similarity without overfitting, ensuring better generalization in zero-shot speech synthesis tasks.
A.3 Prior distribution
The experiments in Table 5 reveal that increasing the prior variance $\sigma^{2}$ in Equation 4 introduces a systematic trade-off between synthesis accuracy and prosodic continuity. For continuation tasks, higher $\sigma^{2}$ values yield improved WER compared to lower settings, suggesting that moderate noise injection enhances the model’s adaptability to contextual variations. This improvement, however, coincides with reduced similarity scores, indicating that while the model becomes more robust to token transitions, excessive variance weakens temporal coherence between adjacent frames. In cross-sentence synthesis, $\sigma^{2}=0.1$ achieves optimal performance. While WER values remain stable across $\sigma^{2}$ settings and similarity metrics degrade sharply at higher $\sigma^{2}$ (e.g., $\sigma^{2}=0.2$). This underscores that cross-sentence continuity relies on tighter prior conditioning (lower $\sigma^{2}$) to preserve acoustic relationships between sentence boundaries.
Table 5: Effect of prior distribution variance ($\sigma^{2}$) on continuation and cross-sentence zero-shot speech synthesis tasks.
$\sigma^{2}$ Continuation Cross-Sentence
WER-C WER-H SIM-r SIM-o WER-C WER-H SIM-r SIM-o
0.10 1.53 2.27 0.539 0.513 2.20 2.89 0.654 0.619
0.15 1.48 2.20 0.524 0.499 2.28 3.07 0.628 0.596
0.20 1.48 2.14 0.502 0.478 2.20 2.91 0.600 0.570

### v2

@@ 5.2. Model Detail
Our model consists of 12 Transformer blocks, designed in line with the architectures of VALL-E (Chen et al., 2025) and MELLE (Meng et al., 2024) to ensure a fair comparison. Each block features 16 attention heads and a feed-forward layer with a dimensionality of 4,096. The decoder incorporates an embedding dimension of 1,024 and ReLU activation function. The input mel-spectrograms are transformed into the model’s embedding space using a three-layer fully connected network. HiFi-GAN vocoder [^1] https://huggingface.co/mechanicalsea/speecht5-tts (Kong et al., 2020) is used for audio reconstruction.
The downsampling operation preserves the even-indexed mel-spectrogram frames as the coarse components. During upsampling, these components are expanded through zero-insertion at odd-indexed positions. Both coarse and fine flow-matching stages share identical backbone architectures comprising three residual blocks, each containing layer normalization, dual fully connected layers, and SiLU activation. The timestep embedding module combines sinusoidal positional encoding with two fully connected layers and SiLU activation. Key architectural differences emerge in the conditioning modules: the coarse stage uses single linear projections for language model outputs, and the fine stage incorporates additional layers to integrate auxiliary features like coarse-mel information.
@@
The model is trained for 2 million iterations on 8 NVIDIA V100 GPUs. Loss coefficients are set to $\beta=0.1$ for the condition loss $\mathcal{L}_{\text{cond}}$ and $\sigma=0.01$ for the stop prediction loss $\mathcal{L}_{\text{stop}}$. The noise variance during training is configured at 0.1. CFG is implemented with dropout probability $p_{\text{drop}}=0.1$. For unconditional setting, we apply a mask of random length between 3 and 10 seconds to speech prompt. During the inference process, we simultaneously process two types of input: one with complete speech prompts and another with masked speech prompts, used as classifier-free guidance. The results are then combined using a weighting factor of $w=1.6$ to produce the final output. The flow-matching framework performs 3 function evaluations using Euler’s method to iteratively generate mel-spectrograms.
5.3. Evaluation Setting
@@
Continuation:
Using the text transcription and the initial 3 seconds of the utterance as a prompt, the model is tasked with seamlessly synthesizing the continuation of the speech.
Cross-sentence:
Given a reference utterance and its transcription from the same speaker as the prompt, along with the text of the target utterance, the model is expected to synthesize the corresponding speech while preserving the speaker’s characteristics.
5.4. Evaluation Metric
Word Error Rate (WER)
: We use two automated speech recognition (ASR) models, the Conformer-Transducer model [^2] https://huggingface.co/nvidia/stt_en_conformer_transducer_xlarge (Gulati et al., 2020) and the HuBERT-Large ASR model [^3] https://huggingface.co/facebook/hubert-large-ls960-ft (Hsu et al., 2021), to evaluate the robustness and intelligibility of the synthesized speech. By contrasting the transcriptions produced by these models with the matching ground truth, the WER is determined. In particular, the WER scores derived from the Conformer-Transducer and HuBERT-Large systems are indicated by WER-C and WER-H, respectively.
Speaker Similarity (SIM)
: WavLM-TDNN [^4] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification (Chen et al., 2022) is employed to extract speaker embeddings from the reference speech and the synthesized speech to assess the in-context learning capability of zero-shot TTS models. The cosine distance, which goes from -1 to 1, is used to measure how similar these embeddings are to one another. There are two assessment indicators taken into account: SIM-r measuring the similarity between synthesized speech and the reconstructed speech prompt, and SIM-o comparing synthesized speech with the original prompt.
Mean Opinion Score (MOS)
: MOS is a widely used metric that reflects the perceived quality of speech, typically rated by listeners on a scale from 1 (bad) to 5 (excellent). With the rapid development of automatic MOS prediction technology (Cooper et al., 2022; Wang et al., 2024, 2023b; Liu et al., 2025), it is now possible to evaluate quality accurately and effectively. We use the RAMP+ model [^5] https://github.com/NKU-HLT/RAMP_MOS (Wang et al., 2023a, 2025b) for speech quality assessment.
Table 1. The predicted MOS results.
System Continuation Cross-Sentence
Ground Truth $4.043_{\pm 0.32}$ $4.043_{\pm 0.32}$
VALL-E $1.828_{\pm 0.24}$ $1.965_{\pm 0.27}$
MELLE $\textbf{3.843}_{\pm 0.38}$ $4.036_{\pm 0.25}$
FELLE $3.836_{\pm 0.39}$ $\textbf{4.157}_{\pm 0.19}$
Table 2. The objective performance comparison of the continuation and cross-sentence zero-shot speech synthesis tasks, with WER-C (%) and WER-H (%) as WER metrics. *The reproduction results are quoted from (Han et al., 2024). Results reported in (Meng et al., 2024) are used. (Meng et al., 2024) provides results of MELLE across various setups, and we show the performance on the same dataset.
@@ 5.4. Evaluation Metric
6. Results
6.1. Comparative Study
This section presents a comprehensive comparison between our proposed approach and existing TTS systems, using predicted MOS scores to assess perceptual quality and objective metrics such as WER and similarity to evaluate linguistic accuracy and fidelity.
From the perspective of word error rate and similarity, we compare FELLE with several speech synthesis models, including ELLA-V, VALL-E R, RALL-E, CLAM-TTS, VALL-E, and MELLE in Table 2. These models are strong baselines in the field of speech synthesis. These models represent a diverse range of approaches, encompassing both discrete representation-based and continuous representation-based methods, providing a valuable benchmark for evaluating our approach. Additionally, the performance of ground truth (GT) is included as a reference point to highlight the upper bounds of these metrics. Another GT, labeled as GT-mel, refers to the reconstruction from the mel-spectrogram with the vocoder.
@@ 6.1. Comparative Study
Table 3. The ablation study from two perspectives: the prior distribution and the generation mechanism. The ✓ denotes methods used in our paper, while others refer to ablation methods. ‘Vanilla Prior’ denotes a standard Gaussian distribution as the prior, ‘HFM’ denotes holistic flow matching that generates complete mel-spectrogram frames through a unified flow-matching process, and ‘DFM’ represents decoupled flow matching where separate flow-matching processes independently generate low- and high-frequency components without cross-band condition.
@@
Figure 3. Continuation performance across different NFEs. Results are averaged over multiple runs and checkpoints.
[…]
Bar chart showing continuation performance at different numbers of function evaluations (NFEs). Each bar represents an average over multiple runs and model checkpoints. Performance varies with the number of NFEs.
6.2. Ablation Study
6.3. Parameter Analysis
In this section, we conduct a detailed analysis of several critical parameters, including the number of function evaluations (NFE), classifier-free guidance scale, the scaling of the C2F-FM network, and prior variance.
@@ 6.3. Parameter Analysis
The impact of the number of function evaluations on TTS performance is shown in Figure 3. In our experiments, we employ the Euler method for numerical integration, where each step corresponds to one function evaluation. Initially, WER decreases as NFE increases, reflecting the improved mapping between prior distributions and acoustic outputs achieved by flow matching. This enhancement leads to better intelligibility and clarity in the generated speech. However, beyond a certain threshold, WER begins to rise, likely due to overfitting or the introduction of excessive distortion caused by additional iterations. In contrast, the similarity metric consistently declines as NFE increases. This suggests that excessive refinement may result in over-smoothing or the loss of fine-grained details, reducing perceptual similarity. These observations highlight a critical trade-off in the use of flow matching: while a moderate number of NFE improves clarity and reduces WER, excessive NFE can degrade both intelligibility and similarity. Therefore, careful tuning of NFE is essential to strike an optimal balance between clarity and similarity.
Figure 4. Continuation performance across CFG scales (1 to 3, step 0.2). CFG scale 1 means no CFG applied.
[…]
Line plot showing continuation performance as CFG scale increases from 1 to 3 in steps of 0.2. The performance generally improves with higher CFG scale, starting from scale 1, which represents no CFG applied.
@@
Table 4. Comparison of performance with the scale of the C2F-FM Network on continuation and cross-sentence zero-shot speech synthesis tasks. The first column indicates the number of residual blocks and the width of fully-connected layers in each block (e.g., 3×512 denotes 3 residual block with 512-dimensional layers). Total parameters (# Params) represent the sum of trainable parameters in both coarse and fine flow-matching networks.
@@
CFG Scale
Scaling of the C2F-FM Network
The results in Table 4 highlight the trends of WER and similarity as the scale of the C2F-FM network increases. The smallest model demonstrates limited learning capacity, resulting in high WER and poor similarity scores, suggesting the limitations of model size on its capabilities. Further scaling to 3×1024 achieves a better balance, with both WER and similarity metrics showing notable improvements. However, when the model size is increased further to 6×1024 and 12×1024, the gains in WER become marginal, and similarity scores begin to degrade, indicating potential overfitting. This suggests that while larger models can capture more complex patterns, excessive scaling may harm generalization, particularly in maintaining high similarity. In conclusion, simply increasing the scale of the C2F-FM network is not sufficient to achieve optimal performance. Future work should focus on developing more efficient architectures and training strategies to improve both WER and similarity without overfitting, ensuring better generalization in zero-shot speech synthesis tasks.
Table 5. Effect of prior distribution variance ($\sigma^{2}$) on zero-shot speech synthesis tasks.
Continuation Task
$\sigma^{2}$ WER-C WER-H SIM-r SIM-o
0.10 1.53 2.27 0.539 0.513
0.15 1.48 2.20 0.524 0.499
0.20 1.48 2.14 0.502 0.478
Cross-Sentence Task
$\sigma^{2}$ WER-C WER-H SIM-r SIM-o
0.10 2.20 2.89 0.654 0.619
0.15 2.28 3.07 0.628 0.596
0.20 2.20 2.91 0.600 0.570
Prior Variance
The experiments in Table 5 reveal that increasing the prior variance $\sigma^{2}$ in Equation 2 introduces a systematic trade-off between synthesis accuracy and prosodic continuity. For continuation tasks, higher $\sigma^{2}$ values yield improved WER, suggesting that moderate noise injection enhances the model’s adaptability to contextual variations. This improvement, however, coincides with reduced similarity scores, indicating that while the model becomes more robust to token transitions, excessive variance weakens temporal coherence between adjacent frames. In cross-sentence synthesis, $\sigma^{2}=0.1$ achieves optimal performance. While WER values remain stable across $\sigma^{2}$ settings and similarity metrics degrade sharply at higher $\sigma^{2}$ (e.g., $\sigma^{2}=0.2$). This underscores that cross-sentence continuity relies on tighter prior conditioning (lower $\sigma^{2}$) to preserve acoustic relationships between sentence boundaries.
7. Conclusion
Acknowledgements.
This work has been supported by the National Key R&D Program of China through grant 2022ZD0116307 and NSF China (Grant No.62271270).